### v1

# Are Neuromorphic Architectures Inherently Privacy-preserving? An Exploratory Study

Abstract.
While machine learning (ML) models are becoming mainstream, including in critical application domains, concerns have been raised about the increasing risk of sensitive data leakage. Various privacy attacks, such as membership inference attacks (MIAs), have been developed to extract data from trained ML models, posing significant risks to data confidentiality. While the mainstream work in the ML community considers traditional Artificial Neural Networks (ANNs) as the default neural model, neuromorphic architectures, such as Spiking Neural Networks (SNNs) have recently emerged as an attractive alternative mainly due to their significantly low power consumption. These architectures process information through discrete events, i.e., spikes, to mimic the functioning of biological neurons in the brain. While the privacy issues have been extensively investigated in the context of traditional ANNs, they remain largely unexplored in neuromorphic architectures, and little work has been dedicated to investigate their privacy-preserving properties.
In this paper, we investigate the question whether SNNs have inherent privacy-preserving advantage. Specifically, we investigate SNNs’ privacy properties through the lens of MIAs across diverse datasets, comparatively with ANNs . We explore the impact of different learning algorithms (surrogate gradient and evolutionary learning), programming frameworks (snnTorch, TENNLab, and LAVA), and various parameters on the resilience of SNNs against MIA. Our experiments reveal that SNNs demonstrate consistently superior privacy preservation compared to ANNs, with evolutionary algorithms further enhancing their resilience. For example, on the CIFAR-10 dataset, SNNs achieve an AUC as low as 0.59 compared to 0.82 for ANNs, and on CIFAR-100, SNNs maintain a low AUC of 0.58, whereas ANNs reach 0.88. Furthermore, we investigate the privacy-utility trade off through Differentially Private Stochastic Gradient Descent (DPSGD) observing that SNNs incur a notably lower accuracy drop than ANNs under equivalent privacy constraints.
journalyear: YYYYjournalvolume: YYYYjournalnumber: Xdoi: XXXXXXX.XXXXXXX
1. Introduction
As ML systems become more sophisticated and widespread, individuals are increasingly relying on these systems, entrusting them with personal and professional data. Consequently, the risk of sensitive information exposure is growing significantly in multiple sectors (Bertino, 2016) including healthcare (Abouelmehdi et al., 2017), finance (Tripathi and Mukhopadhyay, 2020), national security (Thuraisingham, 2002), education (Florea and Florea, 2020) and consumer services (Lee et al., 2015). It is particularly alarming in fields such as healthcare, where the confidentiality of patient data is extremely sensitive as a breach could result in severe personal and financial implications, affecting patient care and institutional credibility (Luo et al., 2018; JM et al., 2018; U.S. Department of Health & Human Services, 2024). In finance, the integrity of financial transactions and records is fundamental for maintaining market stability and preventing fraud (Yu and He, 2021), while in national security , safeguarding classified information is essential to protect national interests and prevent threats to public safety (Kim et al., 2010).
This has led to the development of various privacy attacks targeting ML models to extract sensitive information, including Model Inversion Attacks (Fredrikson et al., 2015), Attribute Inference Attacks (Gong and Liu, 2018), Model Stealing Attacks (Juuti et al., 2019), and Membership Inference Attacks (MIAs) (Shejwalkar et al., 2021). Among these, MIAs are particularly prominent as a significant threat to data privacy (Salomon, 2012), wherein an adversary seeks to ascertain if a specific data point was part of the dataset used to train the model. This intrusion risks exposing sensitive information about individuals in the training dataset, potentially compromising personal data confidentiality (Shokri et al., 2017).
While most research on privacy attacks has concentrated on ANNs, there is limited exploration within neuromorphic computing, particularly with Spiking Neural Networks (SNNs). Designed to emulate the dynamic behavior of biological neurons (Ghosh-Dastidar and Adeli, 2009), SNNs process information through discrete, temporally encoded spikes (Roy et al., 2019), enabling them to handle time-sensitive data efficiently (Hong et al., 2017). Their suitability for edge computing (Shi et al., 2016) and resource-constrained environments further enhances their value, as SNNs effectively process real-world spatiotemporal patterns (Wang et al., 2018). This capability positions SNNs as a promising alternative to traditional neural networks for applications requiring dynamic, real-time data processing.
This work addresses the privacy concerns associated with SNNs through a structured investigation of three core areas: (i) the resilience of ANN and SNN models to Membership Inference Attacks (MIAs), (ii) the factors influencing the privacy-preserving properties of SNNs, and (iii) the privacy-utility trade-off in ANN and SNN models using the DPSGD algorithm.
We consider that the potential resilience of SNNs against MIAs is based on two key aspects. Firstly, the non-differentiable and discontinuous nature of SNNs may weaken the correlation between the model and individual data points, making it more challenging for an attacker to identify the membership of a particular data point in the training set (Meng et al., 2022). Secondly, the unique encoding mechanisms employed by SNNs introduce an additional layer of stochasticity  (Olin-Ammentorp et al., 2021) and variability to the data representation. This added complexity can make it more difficult for an attacker to deduce the unique characteristics of individual data points, thereby making them more indistinguishable.
Investigating the resilience of SNNs against MIAs, our experimental results consistently demonstrated that SNNs exhibit higher resilience to MIAs across the datasets including MNIST, F-MNIST, Iris, Breast Cancer, CIFAR-10, CIFAR-100, and ImageNet. This is evidenced by the lower Area Under the Curve (AUC) values for the Receiver Operating Characteristic (ROC) curves in SNNs compared to their ANN counterparts. Furthermore, our exploration domain encompassed various learning algorithms (surrogate gradient-based and evolutionary learning), programming frameworks (snnTorch, TENNLab, and LAVA), and a wide range of parameters within them, providing a comprehensive analysis of the factors influencing the inherent privacy-preserving properties of SNNs. This in-depth exploration revealed that evolutionary learning algorithms have shown to boost this resilience more effectively compared to the gradient-based methods.
In order to enhance data privacy and explore the compromises between privacy and utility, we study the implementation of the DPSGD algorithm as a privacy defense mechanism (Xu et al., 2021). This introduces controlled noise into the training process, making it harder for attackers to infer the presence of specific data points. However, improved privacy often comes at the cost of reduced model performance, known as the privacy-utility trade-off  (Song et al., 2013). Through the experiments, we observe that SNNs exhibit a notably lower performance drop compared to ANNs for the same level of privacy guarantee. This finding further reinforces our hypothesis regarding the inherent privacy-preserving properties of SNNs.
This paper offers the following notable findings in the field of data privacy, particularly in the context of SNNs:
  • SNNs exhibit higher resilience against MIAs compared to ANNs, with lower AUC scores on CIFAR-10 (SNN: 0.59 vs. ANN: 0.82) and CIFAR-100 (SNN: 0.58 vs. ANN: 0.88), highlighting their potential as a more secure alternative in privacy-sensitive applications.
  • Evolutionary learning algorithms outperform gradient-based methods in MIA resilience, maintaining a consistent AUC of  0.50 across all parameters for Iris and Breast Cancer datasets, compared to 0.57 and 0.55 AUC scores for gradient-based algorithms, respectively.
  • Privacy-utility tradeoff analysis revealing that SNNs incur a lower accuracy drop compared to ANNs when applying DPSGD: For F-MNIST, with privacy guarantees ranging from 0.22 to 2.00, the average accuracy drop is 12.87% for SNNs, significantly lower than the 19.55% drop observed in ANNs.
It should be emphasized that while this investigation highlights SNNs’ enhanced privacy characteristics, these findings are specifically contextualized within privacy-preservation applications. The architectural properties of SNNs that facilitate efficient hardware implementation and reduced computational overhead make them particularly appealing for resource-constrained environments. However, these advantages should not be interpreted as a general superiority of SNNs over ANNs across all applications. Rather, this work is motivated by the intuition that SNNs’ unique information processing mechanisms may offer specific advantages in privacy preservation, warranting systematic investigation in this particular domain.
2. Related Works
As machine learning systems handle increasingly sensitive data, the potential for privacy violations grows. Li et al.  (Liu et al., 2021a) categorize these privacy challenges into two primary areas: privacy attacks and privacy-preserving techniques. Privacy attacks have emerged as a critical concern in ML due to the growing realization that models can act as unintended leak vectors. These attacks can broadly be classified into different types, such as model inversion attacks, model extraction attacks, MIA. Model inversion attacks (Fredrikson et al., 2015) reconstruct input data from model outputs, while model extraction attacks  (Juuti et al., 2019) aim to replicate a model’s functionality without direct access to its architecture or parameters. Among these, MIAs have gained significant attention due to their ability to infer whether a specific data point was used in training a model. According to the survey conducted by  (Hu et al., 2022), MIAs were first proposed in the context of genomic data by Homer et al. (Homer et al., 2008) where an attacker could infer if an individual’s genome was part of a dataset based on summary statistics. Later, Shokri et al.(Shokri et al., 2017) introduced the first systematic MIA framework, showing how adversaries could use shadow models to infer training data membership. Salem et al.  (Salem et al., 2018) reduced the complexity by demonstrating that a single shadow model can perform well compared to using multiple models, and they introduced metric-based attacks that rely on confidence scores and entropy without the need for identical data distribution between shadow and target models. Nasr et al. (Nasr et al., 2019) further expanded MIA into white-box settings, demonstrating that attackers with access to internal model parameters can perform even more effective MIAs. Melis et al. (Melis et al., 2019) extended MIA to federated learning, showing vulnerabilities even in distributed learning settings, where multiple parties collaboratively train a model. Song and Mittal (Song and Mittal, 2021) highlighted the increased privacy risks in generative models such as GANs, where membership inference attacks could be carried out on synthetic data generators. Recent work by Ilyas et al. introduced LiRA (Likelihood Ratio Attack)  (Carlini et al., 2022), a method that further improves the accuracy of MIAs by leveraging confidence scores more effectively to distinguish between training and non-training data points.In 2024, Zarifzadeh et al. (Zarifzadeh et al., 2024) introduced RMIA, a high-power membership inference attack that outperforms prior methods like LiRA and Attack-R, demonstrating superior robustness, particularly at low FPRs, using likelihood ratio tests.
To counteract these privacy attacks, several privacy-preserving techniques have been developed.These techniques range from cryptographic approaches like homomorphic encryption(HE) (Yi et al., 2014) and secure multi-party computation(SMPC) (Zhao et al., 2019) to learning-based defenses such as model obfuscation (Guo et al., 2021) and knowledge distillation (Gou et al., 2021). However, these methods often suffer from computational inefficiencies, particularly in large-scale systems. Another approach is Federated Learning (FL) (Li et al., 2020), which enables collaborative model training without sharing raw data, but remains vulnerable to attacks like MIA. Among these, Differential Privacy (DP) (Dwork, 2006) has gained prominence due to its strong theoretical guarantees and practical applicability in machine learning settings.It provides a robust framework for protecting individual data points by introducing noise during computations. In ML, this concept has been adapted through various algorithms, with DPSG) (Song et al., 2013) being the most prominent which applies DP principles by adding noise to the gradient updates during training, offering a practical way to maintain privacy while training large models without significantly compromising accuracy.
While much of the data privacy research has centered around ANNs, expanding these investigations to SNNs is necessary. SNNs not only offer performance levels comparable to ANNs but also exhibit superior energy efficiency and hardware integration capabilities, positioning them as promising candidates for exploring inherent privacy features. Although privacy attacks on neuromorphic architectures remain underexplored, existing studies have yet to confirm SNNs’ potential resistance to such threats. However, significant strides have been made in privacy-preserving techniques within the neuromorphic domain. For instance, recent efforts by Han et al.  (Han et al., 2023) focus on developing privacy-preserving methods for SNNs, particularly utilizing FL and DP to address both computational efficiency and privacy challenges. Li et al.(Li et al., 2023) introduced a framework that combines Fully Homomorphic Encryption (FHE) with SNNs, enabling encrypted inference while preserving SNNs’ energy efficiency and computational advantages. Similarly, Nikfam et al.(Nikfam et al., 2023) developed an HE framework tailored for SNNs, offering enhanced accuracy over DNNs under encryption, while carefully balancing computational efficiency. Additionally, Safronov et al.(Kim et al., 2022) proposed PrivateSNN, a privacy-preserving framework for SNNs that employs differential privacy to mitigate membership inference attacks, maintaining the energy-efficient nature of SNNs.
3. Background
Figure 1. Schematic representation of the experimental flow. The framework begins by comparing the privacy risks in ANNs and SNNs through MIA(yellow block). The SNN-specific exploration(green block) space evaluates privacy resilience using surrogate gradient algorithms and evolutionary algorithms with varying parameters. Finally, both ANNs and SNNs undergo DPSGD algorithm(blue block) to assess the privacy-utility tradeoff
3.1. Neuromorphic Architecture
Neuromorphic architectures  (Marković et al., 2020) are designed to replicate the neural structures and functionalities of the biological brain, offering an alternative to traditional von Neumann computing systems  (Aspray, 1990). These architectures integrate memory and processing units, enabling massive parallel processing  (Nassi and Callaway, 2009) and event-driven computation  (Hwang and Wu, 2000) .By modeling neurons and synapses that communicate via discrete spikes or events, neuromorphic systems can process data asynchronously, reducing power consumption and latency  (Liu et al., 2021b). Hardware implementations like IBM’s TrueNorth  (Akopyan et al., 2015) and Intel’s Loihi (Davies et al., 2018) demonstrate the potential for scalable, energy-efficient architectures capable of performing complex tasks with a structure that closely resembles biological neural networks.
SNNs are a core component of neuromorphic systems offering a distinct approach compared to traditional ANNs. In SNNs, neurons operate by accumulating charge over time, firing only when a threshold is reached, thereby producing discrete spikes. This introduces a temporal dimension to information processing, as the timing of these spikes encodes key information. SNN models range from basic integrate-and-fire neurons  (Auge et al., 2021), where spikes are triggered by charge accumulation, to more complex, biologically realistic models like the Hodgkin-Huxley model, which captures finer details of neuronal dynamics. This spike-based communication not only mirrors biological neurons but also enables asynchronous data flow, making SNNs particularly suitable for neuromorphic hardware that thrives on event-driven processing. The inclusion of temporal dynamics and synaptic weights in SNNs allows for a more adaptable learning process, which is essential for tasks requiring real-time decision-making and low-latency responses  (Schuman et al., 2022a).
The unique characteristics of SNNs, particularly their spiking-based encoding and asynchronous processing, introduce complexities that may impact their vulnerability to privacy attacks. Traditional ANNs are prone to attacks such as Membership Inference Attacks (MIAs), where adversaries can exploit deterministic output patterns and smooth gradients to infer the presence of specific data points in the training set. In contrast, SNNs operate with discrete, non-differentiable spikes  (Zhang et al., 2018), where the timing and stochastic nature of spike generation add variability to the output. Neurons in SNNs accumulate charge from inputs and fire based on probabilistic thresholds, which can lead to diverse spike patterns even for similar inputs. This inherent variability and the discontinuous nature of SNN activations complicate the tracing of input-output relationships, making it harder for adversaries to identify patterns necessary for MIAs. Although this suggests a potential for increased privacy resilience, especially when combined with techniques like DPSGD, this remains a hypothesis. Our investigation explores whether SNNs, due to their spike-based processing, exhibit lower performance degradation under privacy constraints when compared to ANNs.
3.2. Membership Inference Attack (MIA)
MIA is a privacy attack that enables adversaries to infer the presence of individual data in a machine learning model’s training dataset. This attack exploits the differences in the model’s behavior when queried with data points that were part of the training set versus those that were not  (Rahman et al., 2018). Typically, models tend to exhibit higher confidence or different error patterns for samples they have encountered during training, compared to new unseen data  (Nguyen et al., 2015). By analyzing these patterns, attackers can infer sensitive information about the training data, which can lead to privacy breaches  (De Cristofaro, 2020), especially in applications dealing with confidential or personal data.
MIA involves training attack models or employing statistical tests to distinguish between the model responses on training and non-training data. The success of these attacks largely depends on the model’s overfitting to its training data and the distinctiveness of the model’s responses to individual samples. MIAs not only breach the privacy of data but also expose weaknesses in a model’s ability to generalize (Gomm et al., 2000).
3.3. Differentially Private Stochastic Gradient Descent (DPSGD)
Differential Privacy (DP)  (Dwork, 2006) is a mathematical framework that aims to ensure that the output of a computation remains nearly unchanged irrespective of whether an individual’s data is included in the dataset, thus offering strong privacy guarantees. It provides a framework to quantify the impact on privacy when a single data point in a dataset is changed. The privacy parameters $\epsilon$ (privacy budget) and $\delta$ (delta parameter) control this impact. The core principle is expressed in the inequality:
$P(M(D)\in S)\leq e^{\epsilon}P(M(D^{\prime})\in S)+\delta$
where $D$ and $D^{\prime}$ are datasets differing by one element, $M$ is a randomized algorithm, and $S$ is a subset of potential outputs. This inequality limits the probability of $M$ producing an output within $S$ when applied to $D$ is limited by the exponential of $\epsilon$. This probability is then multiplied by the probability of $M$ producing the same output from $D^{\prime}$. Additionally, it is adjusted by a small term $\delta$. Here, $\epsilon$ modulates the sensitivity of the output to input changes—lower values imply stronger privacy by reducing the allowable change in output probabilities. $\delta$, ideally a small value near zero, accounts for the rare cases where this privacy guarantee might fail, thus ensuring that changes to any single data point in $D$ minimally affect the output, reinforcing data privacy across the dataset.
DPSGD utilizes the principles of DP by incorporating a mechanism of noise addition directly into the gradient descent process used for training machine learning models. In DPSGD, the key step involves perturbing the gradients computed during each training iteration with noise that is calibrated to the sensitivity of the function being optimized. This sensitivity measures how much the output of a function can change in response to changes in its input, which in the context of machine learning, translates to how much a single training example can influence the overall model parameters. The noise added is typically drawn from a Gaussian distribution  (Kassam, 2012), scaled $\epsilon$ and the desired level of privacy guarantee, $\delta$.  (OpenAI, 2024; Anthropic, 2024)
The function of DPSGD can be expressed mathematically as:
$\theta_{t+1}=\theta_{t}-\eta\left(g_{t}+\mathcal{N}(0,\sigma^{2}\mathbf{I})\right)$
where $\theta_{t}$ represents the model parameters at iteration $t$, $\eta$ is the learning rate, $g_{t}$ is the gradient of the loss function with respect to $\theta_{t}$, clipped to a norm bound $C$, and $\mathcal{N}(0,\sigma^{2}\mathbf{I})$ denotes the Gaussian noise added to the gradient, with $\sigma$ being determined by $C$, $\epsilon$, and $\delta$.
4. Methodology
This study aims to assess whether SNNs demonstrate an advantage in privacy resilience compared to ANNs. The methodology is divided into three experimental phases: (1) evaluating privacy risk under MIAs for both ANNs and SNNs, (2) investigating SNN privacy vulnerability across various algorithms, and (3) examining the privacy-utility tradeoff using Differentially Private Stochastic Gradient Descent (DPSGD).
These phases are outlined in the diagram (Figure 1) to illustrate the workflow and interconnections between the models and methods used.
4.1. Comparison of Privacy Vulnerability between ANNs and SNNs
To compare privacy risks, we applied MIAs to both ANNs and SNNs across equivalent architectures: convolutional baseline models, ResNet18, and VGG16. In this phase, shadow models were trained on labeled datasets (Figure 1, top-right) to replicate the behavior of the target model, and attack models were trained to determine the membership of specific data points in the training set. The effectiveness of these attacks was measured using the Area Under the ROC Curve (AUC) for both ANN and SNN models. This comparison allows for a direct analysis of the vulnerability of each model type to MIAs (Figure 1, yellow block).
4.2. Algorithmic Exploration within the SNN Architecture
The second phase focused on evaluating SNNs under different algorithmic conditions to assess the consistency of their privacy resilience. Three frameworks were explored: Surrogate Gradient Algorithm, Evolutionary Algorithm, and Intel’s LAVA framework.
Surrogate Gradient Algorithm: The baseline SNN model was tested with three different encoders (Delta, Latency, Delta Modulation) from the snntorch library (Figure 1, left section, "SNN Exploration Space"). The goal was to analyze the impact of these encoding mechanisms on the model’s susceptibility to MIAs by examining variations in spike generation methods and their effect on attack success rates.
Evolutionary Algorithm: Using the TennLab framework, we systematically varied parameters such as Population Size, Crossover Rate, Mutation Rate, Processor configurations, and Encoder settings. These variations were implemented to determine their impact on the privacy vulnerability of SNN models (Figure 1, green block).
4.3. Privacy-Utility Trade-off Analysis
In the final phase of our experiments, we applied DPSGD to both ANN and SNN models (Figure 1, bottom) to evaluate the privacy-utility tradeoff. DPSGD introduces noise into the gradient updates during training, providing differential privacy protection at the expense of model accuracy. We implemented DPSGD on the convolutional, ResNet18, and VGG16 architectures to quantify the impact of privacy measures on model performance (accuracy) and privacy (attack AUC). This allowed for a comparative analysis of the privacy-utility balance between ANNs and SNNs under identical privacy constraints (Figure 1, blue block).
5. Experimental Framework and Setup
5.1. Dataset and Model Architecture
Table 1. Model Architectures and Configurations
Network Variant Structure Params*
Baseline ANN $\bullet$ 2 Conv (32,64 filters) $\sim$2.3M
(ConvNet) $\bullet$ 2 MaxPool
$\bullet$ 2 FC (1000, num_classes)
$\bullet$ ReLU
SNN $\bullet$ 2 Conv (32,64 filters) $\sim$2.3M
$\bullet$ 2 MaxPool
$\bullet$ LIF neurons
$\bullet$ Temporal processing
ResNet18 ANN $\bullet$ 4 BasicBlocks (64$\to$512) $\sim$11.7M
$\bullet$ GroupNorm
$\bullet$ Skip connections
$\bullet$ Adaptive pool
SNN $\bullet$ 4 BasicBlocks (64$\to$512) $\sim$11.7M
$\bullet$ GroupNorm+BNTT
$\bullet$ Spike residuals
$\bullet$ Membrane threshold
VGG16 ANN $\bullet$ 13 Conv (64$\to$512) $\sim$138M
$\bullet$ 5 MaxPool
$\bullet$ 3 FC (4096, classes)
$\bullet$ GroupNorm, ReLU
SNN $\bullet$ 13 Conv (64$\to$512) $\sim$138M
$\bullet$ 5 AvgPool
$\bullet$ Binary spikes
$\bullet$ Membrane reset
*Params vary with input channels (1/3) and classes (10/100)
We evaluate our method on both image and tabular datasets. For image classification tasks, we use MNIST (LeCun et al., 2010) and Fashion-MNIST (Xiao et al., 2017) (28×28 grayscale images, 10 classes), CIFAR-10 (Recht et al., 2018) and CIFAR-100 (Sharma et al., 2018) (32×32 RGB images, 10 and 100 classes). We also include two tabular datasets: Iris (Omelina et al., 2021) (4 features, 3 classes) and Breast Cancer  (Zwitter and Soklic, 1988) (30 features, 2 classes).
As shown in Table  1, we implement three model architectures in both ANN and SNN configurations. The baseline model uses two convolutional layers for image data and a fully connected architecture for tabular data. ResNet18 consists of four basic blocks with group normalization and skip connections. VGG16 uses 13 convolutional layers with filter sizes increasing from 64 to 512, followed by three fully connected layers. For the SNN variants, we replace ReLU activations with leaky integrate-and-fire (LIF) neurons and add temporal processing components. We implement all ANN models using PyTorch (Ketkar et al., 2021). For SNN implementations, we use snnTorch for the baseline model and SpikingJelly (Fang et al., 2023) for ResNet and VGG16 architectures.
5.2. MIA
Figure 2. Membership Inference Attack(MIA) Framework
We implement the MIA framework using a two-model approach as illustrated in Figure 2. The experimental setup includes a ’target’ model and a ’shadow’ model, where we evaluate three different architectures: baseline model, ResNet18, and VGG16. The shadow model simulates target’s functionality by training on 80% of the same data, maintaining identical architectural framework as the target model. For the attack phase, we employ a Support Vector Machine (SVM) as our attack model, chosen for its effectiveness as a binary classifier to determine whether a data point is a member or non-member of the training set. We evaluate the models by querying with their respective training and test sets, labeling the responses as ’IN’ if the data was in the model’s training set and ’OUT’ otherwise. To prevent the classifier’s bias towards the more frequent classes, we implement undersampling of the predominant class. The SVM attack model, trained on these labeled predictions from the shadow model, provides a quantified measure of the model’s susceptibility to MIA through its ability to accurately classify membership status.
5.3. DPSGD
Figure 3. Differential Privacy via Stochastic Gradient Descent (DPSGD) Algorithm
We implement DPSGD as shown in Figure 3 using the "Privacy Engine" module from Opacus (Yousefpour et al., 2021). The DPSGD algorithm  1 begins by selecting a minibatch $L_{t}$ uniformly from the set of indices $\{1,\dots,N\}$ with size $L$. For each data point $x_{i}$ in the minibatch, the gradient of the loss function $\mathcal{L}(\theta_{t},x_{i})$ with respect to the parameters $\theta_{t}$ is computed as $g_{t}(x_{i})$. These gradients are then clipped using an $L2$ norm operation, $\hat{g}_{t}(x_{i})=g_{t}(x_{i})/\max(1,\|g_{t}(x_{i})\|_{2}/C)$. Gaussian noise $\mathcal{N}(0,\sigma^{2}C^{2}I)$ is added to the average of the clipped gradients, with $\sigma$ being the noise scale, to ensure differential privacy. The noisy gradient $\tilde{g}_{t}$ is used to update the model parameters: $\theta_{t+1}=\theta_{t}-\eta_{t}\tilde{g}_{t}$.
To investigate the privacy-utility tradeoff, we define the same target privacy budget ($\varepsilon$) of range 0.1-2.00, while fixing the privacy compromise $\delta=1e-5$ and evaluate the utility of the SNN architectures comparatively with their ANN counterparts.
Algorithm 1 DPSGD Algorithm
1:Input: Dataset $\{x_{1},\dots,x_{N}\}$, loss function $\mathcal{L}(\theta)=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(\theta,x_{i})$
2:Parameters: Learning rate $\eta$, noise scale $\sigma$, batch size $L$, gradient norm bound $C$
3:Initialize $\theta_{0}$ randomly
4:for $t=1$ to $T$ do
5:     Take a random sample $L_{t}$ with sampling probability $\frac{L}{N}$
6:     for each $i\in L_{t}$ do
7:         Compute gradient $g_{t}(x_{i})\leftarrow\nabla_{\theta}\mathcal{L}(\theta_{t},x_{i})$
8:         Clip gradient $g_{t}(x_{i})\leftarrow\frac{g_{t}(x_{i})}{\max(1,\frac{\|g_{t}(x_{i})\|_{2}}{C})}$
9:     end for
10:     Add noise $\tilde{g}_{t}\leftarrow\frac{1}{L}\left(\sum_{i\in L_{t}}g_{t}(x_{i})+\mathcal{N}(0,\sigma^{2}C^{2}I)\right)$
11:     Update parameters $\theta_{t+1}\leftarrow\theta_{t}-\eta\tilde{g}_{t}$
12:end for
13:Output: Return $\theta_{T}$ and compute privacy cost $(\epsilon,\delta)$
5.4. SNN Exploration Space
In the SNN exploration space as shown in Figure 1, we investigate two major learning algorithms: surrogate gradient-based learning  (Neftci et al., 2019) and evolutionary learning (Schuman et al., 2020b).
5.4.1. Surrogate Gradient based Learning:
Surrogate gradient-based learning algorithms are used to train Spiking Neural Networks (SNNs) by approximating the non-differentiable spiking function with a differentiable surrogate. This allows the application of gradient-based optimization techniques, such as backpropagation, to SNNs. By replacing the true gradient of the spike function with a smooth, differentiable approximation, these algorithms enable the effective training of SNNs. In our experiments, we employ the Arc-tangent  (Rajan et al., 2006) surrogate gradient algorithm through the snnTorch  (SNN-Torch Development, 2021) framework and the Spike-Timing-Dependent Plasticity (STDP)  (Caporale and Dan, 2008) learning rule using the LAVA (Intel Labs, 2023) framework by Intel.
snnTorch Framework:
snnTorch is a deep learning framework specifically designed for SNNs built on PyTorch. For our experiments, we utilized the arc tangent learning algorithm within the framework. This algorithm approximates the non-differentiable spiking function with an arc tangent function. This framework was applied to MNIST, F-MNIST, Iris and Breast Cancer datasets to evaluate the resilience of SNNs against MIA.
In SNNs, encoding refers to the process of converting input data into spike trains that the network can process and it determines how input signals are represented as spikes, which regulates performance and behavior of the SNNs.
We explore the impact of different encoding schemes on privacy preserving properties of SNNs. The three encoding schemes  (Eshraghian, 2023) we investigated on snnTorch are:
  • Rate Encoding: Transforms input features into spikes by representing each feature as a probability of spike occurrence at each time step, with the neuron’s firing rate directly tied to the intensity of the input signal. For our experiments, we set number of steps to 10.
  • Latency Encoding: Encodes information based on the timing of spikes, where inputs with higher values result in earlier spikes, effectively using the temporal dimension to convey input magnitude. We used time step of 10 and set the RC constant, Tau to 5, and Threshold to 0.1 for our experiments.
  • Delta Encoding: Event-driven and produces spikes in response to changes in input features over time, making it adept at capturing dynamic variations in data. In our experiments, we set the threshold to 0.1.
LAVA Framework:
LAVA is an innovative open-source software framework designed for neuromorphic computing  (Roy et al., 2019). It provides a versatile and flexible environment for developing, simulating, and deploying neuromorphic applications. By abstracting hardware details through a process-based model, LAVA enables the creation of scalable and modular systems that can operate asynchronously across various neuromorphic and conventional hardware platforms. The framework supports rapid prototyping and detailed optimization, making it accessible to both researchers and developers interested in leveraging the unique capabilities of neuromorphic technologies. LAVA aligns particularly well with Intel’s Loihi chip  (Davies et al., 2018), a specialized neuromorphic processor. LAVA provides a seamless interface for developing applications that can efficiently run on Loihi. The LAVA framework also supports Spike-Timing-Dependent Plasticity (STDP), a biological learning rule that adjusts synaptic strengths based on the precise timing of spikes. This mechanism allows SNNs to learn temporal patterns and adapt to dynamic inputs. In our experiment, we used framework to evaluate the model’s resilience against privacy attack (MIA).
Figure 4. ROC curves comparing ANN and SNN models across different datasets: (a) MNIST, (b) F-MNIST, (c) CIFAR-10, and (d) CIFAR-100.
Table 2. Comparison of ANN and SNN Resilience to MIA Across Diverse Datasets
Dataset Architecture ANN SNN
Train Acc Test Acc Attack AUC Train Acc Test Acc Attack AUC
MNIST Baseline 99.96% 99.21% 0.59 99.95% 99.22% 0.52
ResNet18 99.98% 99.57% 0.61 99.96% 99.51% 0.51
VGG16 99.67% 99.50% 0.61 99.58% 99.37% 0.51
F-MNIST Baseline 99.52% 92.77% 0.64 99.42% 92.44% 0.54
ResNet18 98.72% 93.66% 0.67 99.12% 91.79% 0.53
VGG16 97.14% 93.06% 0.62 96.53% 90.34% 0.55
CIFAR-10 Baseline 99.24% 73.20% 0.82 99.13% 72.99% 0.59
ResNet18 97.65% 90.81% 0.63 96.74% 86.85% 0.55
VGG16 97.45% 88.74% 0.67 79.77% 74.92% 0.53
CIFAR-100 Baseline 98.31% 42.46% 0.88 99.42% 39.92% 0.58
ResNet18 98.66% 70.92% 0.77 88.43% 60.65% 0.66
VGG16 81.61% 58.32% 0.70 60.53% 52.32% 0.58
Iris Baseline 100% 96.67% 0.77 97.04% 96.67% 0.57
Breast Cancer Baseline 99.54% 97.37% 0.65 99.62% 96.98% 0.55
Figure 5. Overfitting analysis in ANN and SNN models across datasets: F-MNIST, Breast Cancer, and CIFAR-10
5.4.2. Evolutionary Learning Algorithm:
Evolutionary algorithms, such as the Evolutionary Optimization for Neuromorphic Systems (EONS)  (Schuman et al., 2020b), are capable of enabling the rapid prototyping of SNN applications. These applications can be adapted to hardware constraints and various learning scenarios, including classification  (Schuman et al., 2020a, 2022a) and control tasks (Patton et al., 2021; Schuman et al., 2022b). EONS, implemented within the TENNLab framework, facilitates the co-design process between neuromorphic hardware and software, helping developers optimize neural network structures within the constraints of neuromorphic hardware. EONS interacts seamlessly with a wide variety of devices  (Schuman et al., 2020b), architectures  (Schuman et al., 2020c), and application  (Plank et al., 2018) without necessitating any modifications to its underlying algorithm. This approach, implemented within the TENNLab framework, utilizes an evolutionary algorithm to optimize neuromorphic network structure and weights. This process begins with generating a population of potential network solutions, which can be either randomly created or seeded with pre-existing networks  (Schuman et al., 2020b). The EONS process then evaluates these networks for fitness based on user-defined criteria and uses common genetic algorithm selection techniques such as tournament or fitness score to choose the most promising networks for reproduction. During reproduction, selected parent networks undergo crossover to swap segments of their structure and merging, which combines their entire structures into a single offspring. Mutation operations introduce random modifications to a network’s nodes and edges, facilitating diversity and adaptation in the population. These operations are performed on a generalized network representation consisting of nodes and edges, allowing flexibility and adaptability across different hardware implementations. The entire EONS cycle, designed to optimize network parameters and structure efficiently, repeats until achieving desired performance metrics.
In our experiments we depict the impact of different EONs parameters, encoders and neuro processors inside the framework on the vulnerability of the SNN model against MIA on Iris and Breast Cancer dataset.
Population Size: In the EONS approach, population consists of potential solutions that form the genetic pool for evolution. This initial population can be randomly generated or seeded with pre-existing solutions. Each network is evaluated, and the best-performing networks are selected for reproduction. In our experiments, we vary the population size over 50, 100, 200 and 400 to understand its impact on resilience against MIAs.
Mutation Rate: The mutation rate in the EONS approach specifies the frequency at which random changes are introduced to the solutions. These mutations are essential for exploring new solution spaces and preventing the population from stagnating in local optima. Mutations can be structural, such as adding or deleting nodes and edges, or they can involve changes to parameters like thresholds or weights. In our experiments, we vary the mutation rate over 0.1, 0.5 and 0.9 to study its effects on the inherent privacy preservation in SNNs.
Crossover Rate: The crossover rate determines how often parts of two solutions are recombined to create new solutions, enhancing genetic diversity. In EONS, the algorithm uses a node-edge recombination method, where it mixes components from parent networks to form child networks. By varying the crossover rate, the algorithm maintains a diverse genetic pool which is essential for exploring the solution space and identifying high-performing network configurations. By varying the crossover rate over 0.1, 0.5 and 0.9, we aim to explore its impact on the resilience of SNNs against MIAs.
Neuroprocessors: We explore 3 neuroprocessors available in the TENNLab framework:
  • RISP  (Plank et al., 2022): A lightweight neuro processor employing an integrate-and-fire model with discrete time steps for accumulating and evaluating action potentials, suitable for networks with integer synaptic delays.
  • Caspian  (Kulkarni et al., 2021): Provides a high-level API and a fast spiking simulator integrated with FPGA  (Kuon et al., 2008) architecture, enhancing development and deployment of neuromorphic solutions in size, weight, and power-constrained environments.
  • RAVENS  (Foshie et al., 2023): A versatile neuroprocessor from TENNLab with multiple implementations including software simulation, microcontroller, FPGA, ASIC (Good and Benaissa, 2008), and Memristive ASIC  (Weiss, 2022)(mRAVENS), catering to a wide range of computational needs in neuromorphic systems.
Encoding Techniques:
  • Flip Flop Technique: It assigns inverted percentage values in even-numbered bins, ensuring smoother transitions and preserving information about minimum values. It is effective in applications like proximity sensing in LIDAR systems  (Plank et al., 2019).
  • Triangle Technique: It smooths input space by overlapping bins where values rise to 100% at bin boundaries and then fall, facilitating a more gradual representation of input data. This is useful in refined control applications.
Figure 6. ROC curves showing the impact of Rate, Latency, and Delta modulation encoding under MIA for MNIST, F-MNIST, CIFAR-10, Iris, and Breast Cancer datasets.
Figure 7. ROC curves comparing the impact of EONs Parameters in TennLab Framework under MIA on (a)Breast Cancer and (b)Iris Dataset
6. Results
6.1. MIA Assessment
In this section, we evaluate the resilience of both ANN and SNN models to MIAs across multiple architectures, including the baseline, ResNet18, and VGG16 models for each dataset. The evaluation metric used is the Area Under the Curve (AUC) of the Receiver Operating Characteristic (ROC) curve, where lower AUC values signify greater resistance to MIAs.
Our findings consistently demonstrate that SNNs achieve lower AUC values than their ANN counterparts across all datasets, highlighting their stronger inherent resilience against privacy attacks. For instance, on the MNIST dataset (Figure 4a), SNN models yield an AUC as low as 0.51, in contrast to ANNs, which range between 0.59 and 0.61. This pattern persists in more complex datasets like CIFAR-10 (Figure 4c) and CIFAR-100 (Figure 4d), where SNNs achieve respective AUCs of 0.53 and 0.58 across model variations, compared to ANNs which reach AUCs up to 0.82 and 0.88, respectively.
As shown in Table 2, across all architectures, SNNs not only maintain a robust classification accuracy but also exhibit significantly lower vulnerability to MIAs, suggesting an advantage in privacy-preserving properties without compromising performance. The contrast is particularly pronounced in high-dimensional data like CIFAR-100, where ANNs display considerable susceptibility to MIAs (AUC values nearing 0.88), whereas SNNs consistently maintain lower AUCs across all model architectures.
Is SNN resilience driven by ANN overfitting?
To assess whether the improved resilience of SNNs is simply a result of ANN overfitting, we examined both ANN and SNN models trained on the baseline architecture across multiple datasets, including F-MNIST, Breast Cancer, and CIFAR-10. Overfitting is known to increase a model’s vulnerability to MIA by reducing the generalizability of the model. Overfitted models tend to memorize training data, which increases the likelihood that an adversary can successfully infer whether a particular data point was part of the training set, thus leading to a higher MIA AUC.
As demonstrated in Figure 5, both ANN and SNN models exhibit comparable levels of overfitting, as reflected by the similar gaps between their training and testing accuracies. For example, in F-MNIST(Figure 5a), the average overfitting for ANN is 4.13%, while for SNN, it is 4.67%. In the Breast Cancer dataset(Figure 5b), the average overfitting values are much lower, at 1.57% for ANN and 2.63% for SNN. These results indicate that overfitting occurs similarly in both ANNs and SNNs, ruling out overfitting in ANNs as the sole explanation for the enhanced resilience of SNNs against MIAs.
The pronounced overfitting in CIFAR-10 (Figure 5c) can be attributed to the dataset’s complexity and the relatively shallow two-convolution-layer architecture, which, over extended training epochs, struggles to generalize effectively and tends to memorize the training data. Despite these similar overfitting trends, SNNs consistently exhibit lower MIA AUCs than their ANN counterparts. This suggests that the observed resilience of SNNs against MIAs is not due to overfitting in ANNs, but rather reflects inherent differences in how SNNs process and encode information, making them less susceptible to privacy leakage.
6.2. SNN Exploration Space Assessment:
In the SNN domain, we analyze the results by examining different learning algorithms, the frameworks utilized for these algorithms, and the specific parameters within each framework.
6.2.1. Surrogate Gradient Algorithm
In this section we analyze the effects of different encoding schemes within the snnTorch framework and the LAVA framework.
Impact of Encoding Schemes in snnTorch Framework:
Figure 6 illustrates the influence of the three encoding schemes available in snnTorch on the vulnerability of SNNs to MIA. For the MNIST (Figure 6a) and F-MNIST (Figure 6b) datasets, Rate and Delta encoding schemes yield similar AUC values, approximately 0.51 and 0.53, respectively, suggesting a marginal effect on MIA susceptibility. By contrast, the CIFAR-10 dataset (Figure 6c) shows higher AUC values across all encoding types, with values reaching up to 0.58, indicating a slightly elevated vulnerability compared to other datasets. The Iris dataset (Figure 6d) displays more pronounced variation, with Delta encoding achieving a slightly higher AUC of 0.56, suggesting a subtle impact of encoding schemes in smaller datasets. Overall, while no consistent pattern emerges regarding AUC across different encoders, the results consistently fall within a lower AUC range, underscoring the inherent resilience of SNNs towards MIAs across datasets and encoding methods.
Impact of LAVA Framework:
In our experimental setup using the LAVA framework, we employed a three-layer feed-forward neural network architecture featuring Leaky Integrate-and-Fire (LIF) neurons. This network was configured with pre-trained synaptic weights.
As Figure 8 demonstrates, the attack on SNN within the LAVA framework results in an AUC of 0.52, which is about 14.75% lower than its ANN (PyTorch) counterpart.
Figure 8. Impact of MIA on MNIST Dataset in LAVA Framework
Figure 9. Model Performance over increasing Privacy Budget ($\varepsilon$) with DPSGD across MNIST, F-MNIST, CIFAR-10, Iris, and Breast Cancer datasets.
6.2.2. Evolutionary Algorithm:
Figure 7a and Figure 7b respectively depict the results from evaluating Evolutionary Optimization parameters within the TennLab Framework on the Breast Cancer and datasets. For Iris, AUCs range from 0.50 to 0.53 across varying population sizes ( 7b(i)), and remain close to 0.50 with different crossover ( 7b(iii)) and mutation rates ( 7b(ii)), indicating consistent protection. Similarly, the Breast Cancer dataset maintains AUCs near 0.50, with minimal impact from changes in population size ( 7a(i)), crossover rate ( 7a(iii)), or processor type ( 7a(iv)) such as RISP, Caspian, or Ravens.
Compared to surrogate gradient learning (Table 2), where Iris and Breast Cancer datasets scored AUCs of 0.57 and 0.55 respectively, evolutionary algorithms consistently achieve lower AUCs, suggesting an enhanced resilience to MIAs. This consistent resilience is particularly notable, as it indicates that evolutionary algorithms may provide a more robust defense against privacy attacks. Given that evolutionary optimization supports practical hardware implementation within the TennLab framework, achieving a notable privacy gain through this approach could offer substantial benefits for secure neuromorphic applications.
6.3. Privacy-Utility Trade off Assessment
Table 3. Privacy-Utility Trade off Comparison between ANN and SNN
Dataset Privacy Budget Range ANN Avg Accuracy Drop SNN Avg Accuracy Drop
MNIST 0.22 - 2.00 7.89% 6.65%
Fashion-MNIST 0.22 - 2.00 19.55% 12.58%
CIFAR-10 0.22 - 2.00 34.43% 27.87%
Breast Cancer 0.22 - 2.00 6.23% 1.93%
Iris 0.22 - 2.00 29.63% 19.04%
To evaluate the privacy-utility trade-off, we applied Differentially Private Stochastic Gradient Descent (DPSGD) with a fixed privacy budget (epsilon, $\epsilon$) of 0.22-2 across both ANN and SNN models, establishing a consistent basis for comparison. Our results, presented in Figure 9 and Table 3, reveal that SNNs generally experience a smaller drop in classification accuracy compared to ANNs under equivalent privacy constraints. For instance, in the MNIST dataset (Figure 9a), SNN models show an average accuracy drop of 6.65%, while ANNs experience a slightly more pronounced decline of 7.89%. This trend is similarly observed in the F-MNIST dataset (Figure 9b), where SNN accuracy decreases by 12.58% compared to a 19.55% reduction in ANNs. Notably, in CIFAR-10 (Figure 9c), the accuracy drop is 27.87% for SNNs and 34.43% for ANNs, demonstrating that SNNs maintain greater utility even as data complexity increases.
The Breast Cancer dataset (Figure 9e) further illustrates this pattern, with SNNs showing only a 1.93% drop in accuracy compared to a 6.23% decrease in ANNs, reflecting the robustness of SNNs in scenarios with stringent privacy requirements. Similarly, in the Iris dataset (Figure 9d), SNNs exhibit a 19.04% accuracy drop, whereas ANNs experience a 29.63% reduction. These results underscore a consistent advantage for SNNs in balancing privacy and utility, as they retain higher performance levels under equivalent privacy budgets.
7. Conclusion
Our experimental framework encompasses various learning algorithms and programming frameworks, providing a thorough analysis of the factors influencing the privacy preserving properties of SNNs. Our key findings demonstrate that SNNs have a greater ability to withstand MIAs compared to ANNs. Specifically, evolutionary learning algorithms have proven to be more effective than gradient-based methods in strengthening the privacy resilience of SNNs. Additionally, our analysis of the privacy-utility trade off reveals that SNNs show less performance degradation than ANNs when we apply the DPSGD algorithm.
These findings suggest promising directions for privacy-sensitive applications, particularly in scenarios where efficient hardware implementation is required. However, it should be noted that these advantages in privacy preservation and hardware efficiency are specific to certain applications and do not indicate SNNs’ superiority across all domains.
As we move forward, our research endeavors will concentrate on the deployment of our models on Intel’s Loihi neuromorphic chip, enabling the execution of these privacy attacks in a hardware setting. This will provide valuable insights into the practical implementation and real-world performance of SNNs. Concurrently, we intend to expand the scope of our evaluations by incorporating more advanced and sophisticated attack methodologies, subjecting SNNs to increasingly potent threats to further assess their resilience and identify potential vulnerabilities. Moreover, we are actively investigating the integration of differential privacy techniques through evolutionary algorithms within the TennLab framework, aiming to enhance the data privacy protection capabilities of SNNs and explore novel approaches to mitigate privacy risks in neuromorphic computing systems. These ongoing efforts will contribute to a comprehensive understanding of the privacy preserving properties of SNNs and guide the development of robust and secure neuromorphic architectures for privacy-sensitive applications.